OpenStudy (anonymous):

PLEASE HELP I WILL FAN AND MEDAL! Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −1 x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

OpenStudy (anonymous):

To use the substitution method to find the x coordinate of the solution, you'll first want to rearrange both equations so they are in y= form. Can you do this for each of the equations?

OpenStudy (anonymous):

Yes I can.

OpenStudy (jasmineflvs):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1......................(i) 4x + 2y = −1 ..................(ii) Since we have to find x, there y should be eliminated, multiply (i) by 2 4x +2y = 2.....................(iii) subtract (ii) from (iii) 0= 3 ,it cannot be possible Hence there is no solution, because both are parallel..........................Ans There is no x value as there is no solution.

OpenStudy (anonymous):

Ok, once you solve each equation for y, you'll have to equations that are both equal to y, so since they are equal to the same thing they must be equal to each other

OpenStudy (jasmineflvs):

there ya go alex

OpenStudy (anonymous):

@Jasmineflvs The question states to use the substitution method, not elimination.

OpenStudy (jasmineflvs):

the answer is there is no solution